# Strong deflection of massive particles in spherically symmetric spacetimes

### v5

@@ I Introduction
[…]
@@ II Deflection angle of massive particles in static and spherically symmetric spacetimes
[…]
In an asymptotically flat spacetime, as the one we are considering, the constant $E$ represents the energy at infinity per unit rest mass of the particle, while the constant $L$ is the angular momentum per unit rest mass of the particle. From the condition $12\mathcal{L}=-$ 1, that is,
@@
$\frac{d\varphi}{dr}$ =±$=\pm\frac{L}{C(r)}\frac{1}{\sqrt{\dfrac{E^{2}}{A(r)B(r)}-\dfrac{1}{B(r)}\left(1+\dfrac{L^{2}}{C(r)}\right)}}$
=±$=\pm\frac{1}{C(r)}\frac{\sqrt{B(r)}}{\sqrt{\dfrac{1}{L^{2}}\left(\dfrac{E^{2}}{A(r)}-1\right)-\dfrac{1}{C(r)}}}\,.$ (11)
@@
In order to simplify the notation, we will adopt the subscript 00 when a quantity is evaluated at $r_{0}$. Furthermore, we define
@@
[…]
@@
[…]
@@ III Unstable circular orbits of massive particles
$\frac{d}{d\tau}\left[\frac{1}{B(r)}\left(\frac{E^{2}}{A(r)}-\frac{L^{2}}{C(r)}-1\right)\right]\biggr{\rvert}_{r=r_{*}}=0\,.$ (23)
@@
$\frac{C^{\prime}(r_{*})}{C(r_{*})}-\frac{A^{\prime}(r_{*})}{A(r_{*})}\left(1+\frac{A^{2}(r_{*})}{A^{\prime}(r_{*})E^{2}}\right)=0\,.$ (24)
@@
From Eq. (29), we deduce that $E\geq\sqrt{8}/3$, which corresponds to the energy of a particle in the innermost stable circular orbit (in units of $2M$, the radius of the innermost stable circular orbit is equal to 3); according to Eq. (30), $r_{c}$ then varies from $3$ to $3/2$ if the energy $E$ varies from $\sqrt{8}/3$ to infinity. However, as we can easily deduce from Eq. (18), we must impose that $E\geq 1$ in order to allow the particle to escape to $r=\infty$. Thus, the allowed range of $E$ is $[1,\infty)$.
@@ IV Deflection angle in the strong deflection limit
According to Eq. (32), any quantity without the subscript 00 is evaluated at $r$. In line with Refs. Bozza2002 ; Feleppa2024 , we approximate $f(z,r_{0},E)$ to second order in $z$, and define
@@
$\beta\coloneqq\frac{\mathcal{F}_{0}^{2}(1-A_{0})^{2}}{2C_{0}^{2}A_{0}^{\prime 3}}\left[2C_{0}C_{0}^{\prime}A_{0}^{\prime 2}\right.\\ \left.+\hskip 2.84526ptA_{0}^{\prime}A_{0}\left(C_{0}C_{0}^{\prime\prime}-2C_{0}^{\prime 2}\right)\right.\\ \left.\hskip 11.38092pt-\hskip 2.84526ptC_{0}^{2}A_{0}^{\prime\prime}A_{0}\left(\frac{C_{0}^{\prime}}{C_{0}}+2\frac{\mathcal{F}_{0}^{\prime}}{\mathcal{F}_{0}}\right)\right]\\ +\frac{A_{0}(1-A_{0})^{2}}{A_{0}^{\prime 2}}\left(\mathcal{F}_{0}^{\prime 2}+\mathcal{F}_{0}\mathcal{F}_{0}^{\prime\prime}\right).$ (39)
@@
$\frac{C_{0}^{\prime}}{C_{0}}-\frac{A_{0}^{\prime}}{A_{0}}\left(1+\frac{A_{0}^{2}}{A_{0}^{\prime}E^{2}}\right)=0.$ (43)
@@
$\ell\coloneqq\frac{\beta_{c}A_{c}^{\prime 2}\sqrt{C_{c}}}{2\mathcal{F}_{c}A_{c}^{3/2}(1-A_{c})^{2}\left(1-\frac{1}{E^{2}}\right)}.$ (58)
@@ V.2 Reissner-Nordström metric
+$+\left.\left\{3x^{2}\left[6x(2+\log 2)+4+\log 32\right]-\log 2\right\}q^{2}\right\},$ (86)
@@ Appendix A Strong deflection limit coefficients for massive particles in Janis-Newman-Winicour spacetime
[…]
@@
$\mathrm{b}_{R,1}(z,x)=\frac{\left(1-x^{2}\right)\left[3x(3x-4)+12\log\frac{3(1+x)}{2}-5\right]}{24\sqrt{2x^{5}(1+x)}z}+\left\{z\left[3x\left(2-z\right)-z\right]\left[12x(2-z)(z-1)+z(6-5z)\right.\right.\\ \left.\left.+\hskip 2.84526pt9x^{2}z(z-2)\right]\left[5+9x^{2}+6x\left(3+\log\frac{9}{4}\right)-12\log\frac{3}{2}+12(x-1)\log(1+x)\right]\right.\\ \left.\hskip 12.80365pt+\hskip 2.84526pt3z^{2}\left[3x(z-2)+z\right]^{2}\left\{x\left[9x^{2}-3x\left(13+\log 16\right)-29+8\log\frac{6561}{32}\right]+64x\log(1+x)\right.\right.\\ \left.\left.\hskip 27.03003pt+\hskip 2.84526pt12(x-1)^{2}\log 3(x+1)-5-12\log 2\right\}+24x\left[3x(2-z)-z\right]\left\{\left[6x(z-1)(5z^{2}-6z-1)\right.\right.\right.\\ \left.\left.\left.\hskip 38.41139pt+\hskip 2.84526pt10+9x^{2}z-15z+2z^{2}(5z-3)\right]\log\frac{3(1+x)}{2}+(z-1)(z+3xz+2)\left[z+3x(1+z)+5\right]\right\}\right\}\\ \times\left[(1+x)(1+3x)\right]\left\{8\sqrt{3}x\left[3x(z-2)+z\right]^{2}\left[z(1+3x\right]^{2}\sqrt{z^{2}(1+x)\left[3x(2-z)-z\right]}\right\}^{-1}.$ (111)

### v6

@@ I Introduction
In what follows, we set $G=c=1$ and use the signature convention ${,+,+,+}\{$-,+,+,+\}. Moreover, we will work in Schwarzschild coordinates $(t,r,\theta,\varphi)$. Greek indices sum over these coordinates.
@@ II Deflection angle of massive particles in static and spherically symmetric spacetimes
$\frac{\partial\mathcal{L}}{\partial\varphi}-\frac{d}{d\tau}\frac{\partial\mathcal{L}}{\partial\dot{\varphi}}=0\hskip 5.69046pt\Rightarrow\hskip 5.69046pt\dot{\varphi}=\frac{L}{C(r)},\hskip 5.69046ptL=\text{constant}.$ (7)
In an asymptotically flat spacetime, as the one we are considering, the constant $E$ represents the energy at infinity per unit rest mass of the particle, while the constant $L$ is the angular momentum per unit rest mass of the particle. From the condition $2\mathcal{L}=-1$, that is,
@@
$\frac{d\varphi}{dr}$ $=\pm\frac{L}{C(r)}\frac{1}{\sqrt{\dfrac{E^{2}}{A(r)B(r)}-\dfrac{1}{B(r)}\left(1+\dfrac{L^{2}}{C(r)}\right)}}$
$=\pm\frac{1}{C(r)}\frac{\sqrt{B(r)}}{\sqrt{\dfrac{1}{L^{2}}\left(\dfrac{E^{2}}{A(r)}-1\right)-\dfrac{1}{C(r)}}}\,.$ (11)
@@
In order to simplify the notation, we will adopt the subscript 0 when a quantity is evaluated at $r_{0}$. Furthermore, we define
@@
$\Delta\varphi=2\int_{r_{0}}^{\infty}\frac{\sqrt{B}}{\sqrt{C}\sqrt{\frac{C}{C_{0}}\frac{A_{0}}{A}\frac{\mathcal{F}^{2}}{\mathcal{F}^{2}_{0}}-1}}dr,$ (16)
@@
$\hat{\alpha}(r_{0},E)=2\int_{r_{0}}^{\infty}\frac{\sqrt{B}}{\sqrt{C}\sqrt{\frac{C}{C_{0}}\frac{A_{0}}{A}\frac{\mathcal{F}^{2}}{\mathcal{F}^{2}_{0}}-1}}dr-\pi.$ (17)
@@ III Unstable circular orbits of massive particles
$\frac{d}{dr}\left[\frac{1}{B(r)}\left(\frac{E^{2}}{A(r)}-\frac{L^{2}}{C(r)}-1\right)\right]\biggr\rvert_{r=r_{*}}=0\,.$ (23)
@@
$\frac{C^{\prime}(r_{*})}{C(r_{*})}-\frac{A^{\prime}(r_{*})}{A(r_{*})}\left(1-\frac{A(r_{*})}{E^{2}}\right)^{-1}=0\,.$ (24)
@@
From Eq. (29), we deduce that $E\geq\sqrt{8}/3$, which corresponds to the energy of a particle in the innermost stable circular orbit (in units of $2M$, the radius of the innermost stable circular orbit is equal to 3); according to Eq. (30), $r_{c}$ then varies from $3$ to $3/2$ if the energy $E$ varies from $\sqrt{8}/3$ to infinity. However, as we can easily deduce from Eq. (18), we must impose that $E\geq 1$ in order to allow the particle to escape to $r=\infty$. Thus, the allowed range of $E$ is $1,\infty)$.
@@ IV Deflection angle in the strong deflection limit
According to Eq. (32), any quantity without the subscript 0 is evaluated at $r$. In line with Refs. Bozza2002 ; Feleppa2024 , we approximate $f(z,r_{0},E)$ to second order in $z$, and define
@@
$\beta\coloneqq\frac{\mathcal{F}_{0}^{2}(1-A_{0})^{2}}{2C_{0}^{2}A_{0}^{\prime 3}}\left[2C_{0}C_{0}^{\prime}A_{0}^{\prime 2}\right.\\ \left.\hskip-28.45274pt+\hskip 2.84526ptA_{0}^{\prime}A_{0}\left(C_{0}C_{0}^{\prime\prime}-2C_{0}^{\prime 2}\right)\right.\\ \left.\hskip 11.38092pt-\hskip 2.84526ptC_{0}^{2}A_{0}^{\prime\prime}A_{0}\left(\frac{C_{0}^{\prime}}{C_{0}}+2\frac{\mathcal{F}_{0}^{\prime}}{\mathcal{F}_{0}}\right)\right]\\ +\frac{A_{0}(1-A_{0})^{2}}{A_{0}^{\prime 2}}\left(\mathcal{F}_{0}^{\prime 2}+\mathcal{F}_{0}\mathcal{F}_{0}^{\prime\prime}\right).$ (39)
@@
$\frac{C_{0}^{\prime}}{C_{0}}-\frac{A_{0}^{\prime}}{A_{0}}\left(1-\frac{A_{0}}{E^{2}}\right)^{-1}=0.$ (43)
@@
$\ell\coloneqq\frac{\beta_{c}A_{c}^{\prime 2}\sqrt{C_{c}}}{2\mathcal{F}_{c}A_{c}^{3/2}(1-A_{c})^{2}\sqrt{1-1/E^{2}}}.$ (58)
@@ V.2 Reissner-Nordström metric
$\hskip-5.69046pt+\left.\left\{3x^{2}\left[6x(2+\log 2)+4+\log 32\right]-\log 2\right\}q^{2}\right\},$ (86)
@@ Appendix A Strong deflection limit coefficients for massive particles in Janis-Newman-Winicour spacetime
$b_{R}(x,\gamma)=-\hskip 1.42262pt2\sqrt{\frac{1+x}{2x}}\log\left[\frac{\left(\sqrt{3x-1}+\sqrt{6x}\right)^{2}}{24x}\right]+(\gamma-1)\int_{0}^{1}\mathrm{b}_{R,1}(z,x)dz,$ (110)
@@
$\mathrm{b}_{R,1}(z,x)=\frac{\left(1-x^{2}\right)\left[3x(3x-4)+12\log\frac{3(1+x)}{2}-5\right]}{24\sqrt{2x^{5}(1+x)}z}+\left\{z\left[3x\left(2-z\right)-z\right]\left[12x(2-z)(z-1)+z(6-5z)\right.\right.\\ \left.\left.\hskip-24.18501pt+\hskip 2.84526pt9x^{2}z(z-2)\right]\left[5+9x^{2}+6x\left(3+\log\frac{9}{4}\right)-12\log\frac{3}{2}+12(x-1)\log(1+x)\right]\right.\\ \left.\hskip 12.80365pt+\hskip 2.84526pt3z^{2}\left[3x(z-2)+z\right]^{2}\left\{x\left[9x^{2}-3x\left(13+\log 16\right)-29+8\log\frac{6561}{32}\right]+64x\log(1+x)\right.\right.\\ \left.\left.\hskip 27.03003pt+\hskip 2.84526pt12(x-1)^{2}\log 3(x+1)-5-12\log 2\right\}+24x\left[3x(2-z)-z\right]\left\{\left[6x(z-1)(5z^{2}-6z-1)\right.\right.\right.\\ \left.\left.\left.\hskip 38.41139pt+\hskip 2.84526pt10+9x^{2}z-15z+2z^{2}(5z-3)\right]\log\frac{3(1+x)}{2}+(z-1)(z+3xz+2)\left[z+3x(1+z)+5\right]\right\}\right\}\\ \times\left[(1+x)(1+3x)\right]\left\{8\sqrt{3}x\left[3x(z-2)+z\right]^{2}\left[z(1+3x\right]^{2}\sqrt{z^{2}(1+x)\left[3x(2-z)-z\right]}\right\}^{-1}.$ (111)